Brendan is going on a candy binge and needs to stock up on sweets. He buys 2 more bags of M&M's as he does Snickers bars, and twice as many bags of skittles as Snickers bars. If a bag of skittles is $1.50, M&M's are $1.25, a Snicker bar costs $1.00, and Brendan spent a total of $34.00, how many bags of skittles did he buy?
Answer by TimothyLamb(4379)

---
x = m&m
y = snickers
z = skittles
---
x = y + 2
z = 2y
125x + 100y + 150z = 3400
---
put the system of linear equations into standard form
---
x - y = 2
2y - z = 0
125x + 100y + 150z = 3400
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = m&m = 8
y = snickers = 6
z = skittles = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php